Shift report 20030302 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Mar 02 2003 - 16:07:55 EST

  • Next message: Apache: "Shift report 20030303 16:00-24:00"
    ------------------------------------------------------------
    Date: Sun Mar  2 16:07:54 2003, Shift: 08:00-16:00
    Supervisor: FR, with: HB
    Log book#14 pages 43 - 43
    
    Shift Summary:
    ------------------------------------------------------------
    
    1) Beam continued from previous shift. Around 11:30 we voted for dump with PHENIX (because of low ZCC rate). A new store started at 14:20 but was lost after about 1 hour because of power supply/magnet problem.
    
    2) Data runs
    - Run 8171 from previous shift: FS at 4deg, 1/4A - MRS at 40deg, 1050B - 400K trig6 events.
    - Run 8172: FS at 4deg, 1/2A - MRS at 40deg, 1050B - 150k trigger 6 events.
    - Run 8173: FS at 4deg, 1/2A - MRS at 40deg, 1050B - 41k trigger 2 events.
    - Runs 8174-8176 are calibration runs.
    - Run 8177: same setting as for Run 8173 - 18k trigger 2.
    
    3) No detector problems observed.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Mar 02 2003 - 16:12:59 EST